# Mutual position of two smooth quadrics over finite fields

### v1

@@ Abstract.
Given two irreducible conics $C$ and $D$ over a finite field $\mathbb{F}_{q}$ with $q$ odd, we show that there are $q^{2}/4+O(q^{3/2})$ points $P$ in $\mathbb{P}^{2}(\mathbb{F}_{q})$ such that $P$ is external to $C$ and internal to $D$. This answers a question of Korchmáros. We also prove the analogous result for higher-dimensional smooth quadric hypersurfaces over $\mathbb{P}^{n-1}$ with $n$ odd, where the answer is $q^{n-1}/4+O(q^{n-\frac{3}{2}})$.
@@ 1. Introduction
For the present paper, we focus on the problem of counting special points relative to two quadric hypersurfaces (defined by two quadratic forms in $n$ variables) viewed as subsets of $\mathbb{P}^{n-1}$. The following paragraph describes the planar case $n=3$ as a motivation.
Let $C\subset\mathbb{P}^{2}$ be a smooth conic defined over a finite field $\mathbb{F}_{q}$. Given a point $P\in\mathbb{P}^{2}(\mathbb{F}_{q})$, there are exactly three possibilities:
@@
Korchmáros [13, Problem 6.2] asked the following natural question.
@@
The prediction in the theorem is intuitive: the “probability” for a point $P$ to be internal or external should be equally likely, leading to $\frac{1}{4}$ chances that $P$ is external to $C$ but internal to $D$. Since the number of points in $\mathbb{P}^{2}(\mathbb{F}_{q})$ is $q^{2}+q+1$, the leading term $\frac{q^{2}}{4}$ in the theorem is reasonable. In the analysis, we should avoid points $P$ which lie on either of the two conics (as the definitions of internal/external do not apply to such points); however, the number of such points is at most $2(q+1)$, which is absorbed by the error term $O(q^{3/2})$. In the next paragraph, we will consider the higher-dimensional version of this question.
@@
Let $q$ be an odd prime power and $n\geq 3$ be odd. Let $C$ and $D$ be two distinct smooth quadrics in $\mathbb{P}^{n-1}$ defined over $\mathbb{F}_{q}$, respectively. The number of points $P$ in $\mathbb{P}^{n-1}(\mathbb{F}_{q})$ external to $C$ but internal to $D$ is
$\frac{q^{n-1}}{4}+O(q^{n-\frac{3}{2}}),$
where the implicit constant in the error term depends only on $n$.
@@
Following the same approach in the proof of Theorem 1.3, we can also prove the same estimate on the number of points $P$ in $\mathbb{P}^{n-1}(\mathbb{F}_{q})$,
  1. (i)
    internal to $C$ and internal to $D$, or
  2. (ii)
    external to $C$ and external to $D$.
Remark 1.5.
@@
Structure of the paper. In Section 2, we introduce the concepts of internal and external points for a given smooth quadric. In Section 3, we borrow tools from character sum estimates to develop a key technical lemma. We prove Theorem 1.3 in Section 4.
@@ 2. Internal and External Points, and Dual quadrics
In this section, we define the concept of an internal and external point to a given quadric hypersurface in $\mathbb{P}^{n-1}$ when $n$ is odd. After computing the discriminant associated with a hyperplane section of a quadric, we give an algebraic criterion for detecting internal/external points. We also explain the relevant notations, recall the classical constructions, and provide further context for our paper. Throughout, $q$ is an odd prime power.
2.1. Internal and external for higher-dimensional quadrics
While the concept of internal and external points has been studied extensively for plane conics, the generalization to higher-dimensional quadrics has received little attention (see Remark 2.5). To motivate our definition, we recall the construction of the dual variety from algebraic geometry. Let $X=\{F=0\}$ be a smooth quadric hypersurface of dimension $n-2$ in $\mathbb{P}^{n-1}$. We use $(\mathbb{P}^{n-1})^{\ast}$ to denote the parameter space of all hyperplanes in $\mathbb{P}^{n-1}$. Associated with $X$, we have a dual hypersurface,
@@ 2.1. Internal and external for higher-dimensional quadrics
It turns out that $X^{\ast}$ is also a smooth quadric hypersurface. Given an $\mathbb{F}_{q}$-point $P\in\mathbb{P}^{n-1}\setminus X$, we can consider the hyperplane $H^{\ast}_{P}\subset(\mathbb{P}^{n-1})^{\ast}$ in the dual space parametrizing those hyperplanes in $\mathbb{P}^{n-1}$ passing through $P$. Observe that the intersection
@@
The words elliptic and hyperbolic appear in finite geometry, namely in the theory of polar spaces [4, Section 4.2], or quadratic spaces over finite fields [20, Section 2]. More precisely, an $(m-1)$-dimensional projective space over $\mathbb{F}_{q}$ equipped with a non-degenerate quadratic form in $m$ variables is isometrically isomorphic to a space equipped with one of the following quadratic forms (see [4, Theorem 3.28] or [20, equation (1.2)]):
@@
The definition of internal and external points for the special case $n=3$ (the planar case) from the introduction is consistent with Definition 2.1. Indeed, the hyperbolic binary form $x_{1}x_{2}$ has two zeros $[0:1]$ and $[1:0]$ in $\mathbb{P}^{1}(\mathbb{F}_{q})$, corresponding to the two tangent lines to the conic passing through $P$. The elliptic binary form $x_{1}^{2}-\lambda x_{2}^{2}$ for a non-square $\lambda\in\mathbb{F}_{q}$ has no zeros in $\mathbb{P}^{1}(\mathbb{F}_{q})$, indicating the absence of tangent lines passing through $P$.
As we saw previously, the presence or absence of solutions dictates whether the point is external or internal. As a measure of how many solutions there are, we use the following concept. The rank of a polar space is the dimension of a maximum totally isotropic subspace. For a polar space equipped with a non-degenerate quadratic form with $2k$ variables, the rank is $k$ in the hyperbolic case, and the rank is $k-1$ in the elliptic case. There is only the parabolic case for a polar space where the number of variables is $12k+$ 1, where the rank is $k+1$ [4, Table 4.1]. Since hyperbolic spaces have larger rank than elliptic ones, the quadratic form in the hyperbolic case has more zeros over $\mathbb{F}_{q}$ compared to the elliptic case (by applying induction on the dimension of the space; see, for example, [15]), justifying Definition 2.1.
When $n$ is even, the polar space $Y_{P}$ is equipped with a quadratic form with $m=n-1$ variables. Thus, $Y_{P}$ is parabolic for every $P\notin X$, and it is unclear how to classify the $\mathbb{F}_{q}$-points of $\mathbb{P}^{n-1}\setminus X$ into the two categories such as internal and external. Indeed, all the quadratic forms associated with a parabolic polar space have the same number of $\mathbb{F}_{q}$-points.
When $n$ is odd, to distinguish the two cases in Definition 2.1, it suffices to study the discriminant of the quadratic form $G_{P}$. The discriminant of a non-degenerate quadratic form $\sum_{1\leq i\leq j\leq m}a_{ij}x_{i}x_{j}$ in $m$ variables is defined by the determinant of the matrix $(a_{ij})$ where $a_{ji}=a_{ij}$ for $i<j$. For the associated quadratic hypersurface, the discriminant is well-defined only inside the quotient $\mathbb{F}^{\ast}_{q}/(\mathbb{F}^{\ast}_{q})^{2}$, where $m$ is necessarily even; thus, we view the discriminant as an element in $\mathbb{F}^{\ast}_{q}/(\mathbb{F}^{\ast}_{q})^{2}$. Lemma 2.4 allows us to detect whether the form is hyperbolic or elliptic based on the value of the discriminant. The next subsection delves further into the specific properties of the discriminant.
@@ 2.2. Discriminant of the hyperplane section of a quadric
Since $X=\{F=0\}$ is smooth, $F$ is a non-degenerate quadratic form. After an invertible linear change of variables, we may assume that $F(\mathbf{x})=a_{0}x_{0}^{2}+a_{1}x_{1}^{2}+\cdots+a_{n-1}x_{n-1}^{2}$ [20, equation (1.1)] with $a_{i}\neq 0$ for each $10\leq i\leq n-$ 1. Note that the discriminant of a quadratic form is invariant under such a change of variables. Let $A$ be the diagonal matrix $(a_{i})$; then we can write $F(\mathbf{x})=\mathbf{x}A\mathbf{x}^{t}$ where $\mathbf{x}$ is a row vector.
@@
Now Lemma 2.2 immediately follows from the lemma below.
@@
Let $a_{0},a_{1},\ldots,a_{n}\in\mathbb{F}_{q}^{*}$. Let $c_{0}=1$, and let $c_{1},c_{2},\ldots,c_{n}\in\mathbb{F}_{q}$. Let $A_{n}$ be the diagonal matrix with diagonal entry $(a_{i}^{-1})_{i=0}^{n}$,
$B_{n}=\begin{pmatrix}\mathbf{c}\\ I_{n}\end{pmatrix}=\begin{pmatrix}c_{1}&c_{2}&\cdots&c_{n}\\ 1&0&\cdots&0\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\end{pmatrix},$
@@
$\det(E_{n})=a_{0}a_{1}\cdots a_{n}\cdot\sum_{i=0}^{n}\frac{c_{i}^{2}}{a_{i}}.$
@@
$M_{n}=B_{n}^{t}A_{n}^{-1}=\begin{pmatrix}a_{0}c_{1}&a_{1}&0&\cdots&0\\ a_{0}c_{2}&0&a_{2}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{0}c_{n}&0&0&\cdots&a_{n}\\ \end{pmatrix}.$
We apply the Cauchy-Binet formula to calculate the determinant of the matrix $E_{n}=M_{n}B_{n}$. Since $M_{n}$ is an $n\times(n+1)$ matrix and $B_{n}$ is an $(n+1)\times n$ matrix, we have $\binom{n+1}{n}=n+1$ summands in the Cauchy-Binet formula. More precisely,
$\det(E_{n})=\sum_{i=0}^{n}\det(M^{(i)}_{n})\cdot\det(B^{(i)}_{n}),$
where $M^{(i)}_{n}$ and $B^{(i)}_{n}$ are the matrices obtained from $M_{n}$ and $B_{n}$ by removing the $(i+1)$-th column and $(i+1)$-th row, respectively. For $i=0$, we have $M^{(0)}_{n}=\operatorname{diag}(a_{1},...,a_{n})$ and $B^{(0)}_{n}=I_{n}$, so
$\det(M^{(0)}_{n})\cdot\det(B^{(0)}_{n})=a_{1}a_{2}\cdots a_{n}=(a_{0}a_{1}...a_{n})\cdot\frac{c_{0}^{2}}{a_{0}}$
@@
$M^{(1)}_{n}=\begin{pmatrix}a_{0}c_{1}&0&\cdots&0\\ a_{0}c_{2}&a_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ a_{0}c_{n}&0&\cdots&a_{n}\\ \end{pmatrix}\quad\quad\quad B^{(1)}_{n}=\begin{pmatrix}c_{1}&c_{2}&\cdots&c_{n}\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\\ \end{pmatrix}$
Hence, $\det(M^{(1)}_{n})\det(B^{(1)}_{n})=(a_{0}a_{2}\cdots a_{n}c_{1})\cdot(c_{1})=(a_{0}a_{1}\cdots a_{n})\cdot\frac{c_{1}^{2}}{a_{1}}$. Similar computation for $i>1$ shows that $\det(M^{(i)}_{n})\det(B^{(i)}_{n})=(a_{0}a_{1}\cdots a_{n})\frac{c^{2}}{a_{i}}$. Thus,
$\det(E_{n})=\sum_{i=0}^{n}\det(M^{(i)}_{n})\cdot\det(B^{(i)}_{n})=a_{0}a_{1}\cdots a_{n}\cdot\sum_{i=0}^{n}\frac{c_{i}^{2}}{a_{i}}$

### v2

@@ Abstract.
Given two irreducible conics $C$ and $D$ over a finite field $\mathbb{F}_{q}$ with $q$ odd, we show that there are $q^{2}/4+O(q^{3/2})$ points $P$ in $\mathbb{P}^{2}(\mathbb{F}_{q})$ such that $P$ is external to $C$ and internal to $D$. This answers a question of Korchmáros. We also prove the analogous result for higher-dimensional smooth quadric hypersurfaces in $\mathbb{P}^{n-1}$ with $n$ odd, where the answer is $q^{n-1}/4+O(q^{n-\frac{3}{2}})$.
@@ 1. Introduction
In the present paper, we focus on the problem of counting special points relative to two quadric hypersurfaces (defined by two quadratic forms in $n$ variables) viewed as subsets of $\mathbb{P}^{n-1}$. The following paragraph describes the planar case $n=3$ as a motivation.
Let $q$ be an odd prime power. Let $C\subset\mathbb{P}^{2}$ be a smooth conic defined over a finite field $\mathbb{F}_{q}$. Given a point $P\in\mathbb{P}^{2}(\mathbb{F}_{q})$, there are exactly three possibilities:
@@
Korchmáros [15, Problem 6.2] asked the following natural question.
@@
The prediction in the theorem is intuitive: the “probability” that a point $P$ is internal or external should be equally likely, leading to a $\frac{1}{4}$ chance that $P$ is external to $C$ but internal to $D$. Since the number of points in $\mathbb{P}^{2}(\mathbb{F}_{q})$ is $q^{2}+q+1$, the leading term $\frac{q^{2}}{4}$ in the theorem is reasonable. In the analysis, we should avoid points $P$ that lie on either of the two conics (as the definitions of internal/external do not apply to such points); however, the number of such points is at most $2(q+1)$, which is absorbed by the error term $O(q^{3/2})$. In the next paragraph, we consider the higher-dimensional version of this question.
@@
$S_{1}$ =$=\{P\in\mathbb{P}^{n-1}(\mathbb{F}_{q})\ |\ P\text{ is internal to }C\text{ and internal to }D\},$
$S_{2}$ =$=\{P\in\mathbb{P}^{n-1}(\mathbb{F}_{q})\ |\ P\text{ is external to }C\text{ and external to }D\},$
$S_{3}$ =$=\{P\in\mathbb{P}^{n-1}(\mathbb{F}_{q})\ |\ P\text{ is internal to }C\text{ and external to }D\},$
$S_{4}$ =$=\{P\in\mathbb{P}^{n-1}(\mathbb{F}_{q})\ |\ P\text{ is external to }C\text{ and internal to }D\}.$
Our main result demonstrates that each of these four sets has size roughly equal to $\frac{q^{n-1}}{4}$.
@@
Let $q$ be an odd prime power and $n\geq 3$ be odd. Let $C$ and $D$ be two distinct smooth quadrics in $\mathbb{P}^{n-1}$ defined over $\mathbb{F}_{q}$, respectively. Then
$|S_{i}|=\frac{q^{n-1}}{4}+O(q^{n-\frac{3}{2}})$
for each $1\leq i\leq 4$, where the implicit constant in the error term depends only on $n$.
@@
Structure of the paper. In Section 2, we introduce the concepts of internal and external points for a smooth quadric. In Section 3.1, we borrow tools from character sum estimates to develop a key technical lemma. We prove Theorem 1.3 in Section 3.2.
@@ 2. Internal and External Points, and Dual quadrics
In this section, we define the concept of an internal and external point to a given quadric hypersurface in $\mathbb{P}^{n-1}$ when $n$ is odd. After computing the discriminant associated with a hyperplane section of a quadric, we give an algebraic criterion for detecting internal/external points. We also explain the relevant notations, recall the classical constructions, and provide further context for our paper. Throughout the section, $q$ is an odd prime power.
2.1. Internal and external points for higher-dimensional quadrics
The concept of internal and external points has been studied extensively for plane conics. Generalizations to higher-dimensional quadrics also appear in the literature [12, Section 1.9]; see also Remark 2.5 for additional references. To motivate the standard definition, we recall the construction of the dual variety from algebraic geometry. Let $X=\{F=0\}$ be a smooth quadric hypersurface of dimension $n-2$ in $\mathbb{P}^{n-1}$. We denote by $(\mathbb{P}^{n-1})^{\ast}$ the parameter space of all hyperplanes in $\mathbb{P}^{n-1}$. Associated with $X$, we have a dual hypersurface,
@@ 2.1. Internal and external points for higher-dimensional quadrics
It turns out that $X^{\ast}$ is also a smooth quadric hypersurface. Given an $\mathbb{F}_{q}$-point $P\in\mathbb{P}^{n-1}\setminus X$, consider the hyperplane $H^{\ast}_{P}\subset(\mathbb{P}^{n-1})^{\ast}$ in the dual space parametrizing those hyperplanes in $\mathbb{P}^{n-1}$ passing through $P$. Observe that the intersection
@@
The words elliptic and hyperbolic appear in finite geometry, namely in the theory of polar spaces [4, Section 4.2], or quadratic spaces over finite fields [22, Section 2]. More precisely, an $(m-1)$-dimensional projective space over $\mathbb{F}_{q}$ equipped with a non-degenerate quadratic form in $m$ variables is isometrically isomorphic to a space equipped with one of the following quadratic forms (see [4, Theorem 3.28] or [22, equation (1.2)]):
@@
The definition of internal and external points in the planar case $n=3$, as given in the introduction, is consistent with Definition 2.1. Indeed, the hyperbolic binary form $x_{1}x_{2}$ has two zeros $[0:1]$ and $[1:0]$ in $\mathbb{P}^{1}(\mathbb{F}_{q})$, corresponding to the two tangent lines to the conic passing through $P$. The elliptic binary form $x_{1}^{2}-\lambda x_{2}^{2}$ for a non-square $\lambda\in\mathbb{F}_{q}$ has no zeros in $\mathbb{P}^{1}(\mathbb{F}_{q})$, indicating the absence of tangent lines passing through $P$.
As we saw previously, the presence or absence of solutions dictates whether the point is external or internal. To quantify the number of solutions, we use the following concept. The rank of a polar space is the dimension of a maximum totally isotropic subspace. For a polar space equipped with a non-degenerate quadratic form with $2k$ variables, the rank is $k$ in the hyperbolic case and $k-1$ in the elliptic case. When the number of variables is $12k+$ 1, the only possibility is the parabolic case, where the rank is $k$ [4, Table 4.1]. Since hyperbolic spaces have larger rank than elliptic ones, the quadratic form in the hyperbolic case has more zeros over $\mathbb{F}_{q}$ compared to the elliptic case (by applying induction on the dimension of the space; see, for example, [17]), justifying Definition 2.1.
When $n$ is even, the polar space $Y_{P}$ is equipped with a quadratic form with $m=n-1$ variables. Thus, $Y_{P}$ is parabolic for every $P\notin X$, and it is unclear how to classify the $\mathbb{F}_{q}$-points of $\mathbb{P}^{n-1}\setminus X$ into the two categories such as internal and external. Indeed, all quadratic forms associated with a parabolic polar space have the same number of $\mathbb{F}_{q}$-points.
When $n$ is odd, to distinguish the two cases in Definition 2.1, it suffices to study the discriminant of the quadratic form $G_{P}$. The discriminant of a non-degenerate quadratic form $\sum_{1\leq i,j\leq m}a_{ij}x_{i}x_{j}$ in $m$ variables is defined by the determinant of the matrix $(a_{ij})$ assuming that $a_{ji}=a_{ij}$ for $i\neq j$. For the associated quadratic hypersurface, the discriminant is well-defined only inside the quotient $\mathbb{F}^{\ast}_{q}/(\mathbb{F}^{\ast}_{q})^{2}$, where $m$ is necessarily even; thus, we view the discriminant as an element in $\mathbb{F}^{\ast}_{q}/(\mathbb{F}^{\ast}_{q})^{2}$. Lemma 2.4 allows us to detect whether the form is hyperbolic or elliptic based on the value of the discriminant. The next subsection delves further into the specific properties of the discriminant.
@@ 2.2. Discriminant of the hyperplane section of a quadric
Since $X=\{F=0\}$ is smooth, $F$ is a non-degenerate quadratic form. After an invertible linear change of variables, we may assume that $F(\mathbf{x})=\alpha_{0}x_{0}^{2}+\alpha_{1}x_{1}^{2}+\cdots+\alpha_{n-1}x_{n-1}^{2}$ [22, equation (1.1)] with $\alpha_{i}\neq 0$ for each $10\leq i\leq n-$ 1. Note that the discriminant of a quadratic form is invariant under such a change of variables. Let $A$ be the diagonal matrix $(\alpha_{i})$; then we can write $F(\mathbf{x})=\mathbf{x}A\mathbf{x}^{t}$ where $\mathbf{x}$ is a row vector.
@@
Now, Lemma 2.2 immediately follows from the lemma below by setting $a_{i}=\alpha_{i}^{-1}$ for each $10\leq i\leq n-$ 1.
@@
Let $a_{0},a_{1},\ldots,a_{n-1}\in\mathbb{F}_{q}^{*}$. Let $c_{0}=1$, and let $c_{1},c_{2},\ldots,c_{n-1}\in\mathbb{F}_{q}$. Let $A_{n}$ be the diagonal matrix with diagonal entry $(a_{i}^{-1})_{i=0}^{n-1}$,
$B_{n}=\begin{pmatrix}\mathbf{c}\\ I_{n}\end{pmatrix}=\begin{pmatrix}c_{1}&c_{2}&\cdots&c_{n-1}\\ 1&0&\cdots&0\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\end{pmatrix},$
@@
$\det(E_{n})=a_{0}a_{1}\cdots a_{n-1}\cdot\sum_{i=0}^{n-1}\frac{c_{i}^{2}}{a_{i}}.$
@@
$M_{n}=B_{n}^{t}A_{n}^{-1}=\begin{pmatrix}a_{0}c_{1}&a_{1}&0&\cdots&0\\ a_{0}c_{2}&0&a_{2}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{0}c_{n-1}&0&0&\cdots&a_{n-1}\\ \end{pmatrix}.$
We apply the Cauchy-Binet formula to calculate the determinant of the matrix $E_{n}=M_{n}B_{n}$. Since $M_{n}$ is an $(n-1)\times n$ matrix and $B_{n}$ is an $n\times(n-1)$ matrix, we have $\binom{n}{n-1}=n$ summands in the Cauchy-Binet formula. More precisely,
$\det(E_{n})=\sum_{i=0}^{n-1}\det(M^{(i)}_{n})\cdot\det(B^{(i)}_{n}),$
where $M^{(i)}_{n}$ and $B^{(i)}_{n}$ are the matrices obtained from $M_{n}$ and $B_{n}$ by removing the $(i+1)$-th column and $(i+1)$-th row, respectively. For $i=0$, we have $M^{(0)}_{n}=\operatorname{diag}(a_{1},...,a_{n-1})$ and $B^{(0)}_{n}=I_{n}$, so
$\det(M^{(0)}_{n})\cdot\det(B^{(0)}_{n})=a_{1}a_{2}\cdots a_{n-1}=(a_{0}a_{1}...a_{n-1})\cdot\frac{c_{0}^{2}}{a_{0}}$
@@
$M^{(1)}_{n}=\begin{pmatrix}a_{0}c_{1}&0&\cdots&0\\ a_{0}c_{2}&a_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ a_{0}c_{n-1}&0&\cdots&a_{n-1}\\ \end{pmatrix}\quad\text{ and }\quad B^{(1)}_{n}=\begin{pmatrix}c_{1}&c_{2}&\cdots&c_{n-1}\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\\ \end{pmatrix}.$
Hence, $\det(M^{(1)}_{n})\det(B^{(1)}_{n})=(a_{0}a_{2}\cdots a_{n-1}c_{1})\cdot(c_{1})=(a_{0}a_{1}\cdots a_{n-1})\cdot\frac{c_{1}^{2}}{a_{1}}$. Similar computation for $i>1$ shows that $\det(M^{(i)}_{n})\det(B^{(i)}_{n})=(a_{0}a_{1}\cdots a_{n-1})\frac{c^{2}}{a_{i}}$. Thus,
$\det(E_{n})=\sum_{i=0}^{n-1}\det(M^{(i)}_{n})\cdot\det(B^{(i)}_{n})=a_{0}a_{1}\cdots a_{n-1}\cdot\sum_{i=0}^{n-1}\frac{c_{i}^{2}}{a_{i}}$